# Uncertainty-Aware 4D Gaussian Splatting for Monocular Occluded Human Rendering

### v1

@@ Abstract
High-fidelity rendering of dynamic humans from monocular videos typically degrades catastrophically under occlusions. Existing solutions incorporate external priors—either hallucinating missing content via generative models, which induces severe temporal flickering, or imposing rigid geometric heuristics that fail to capture diverse appearances. To this end, we reformulate the task as a Maximum A Posteriori estimation problem under heteroscedastic observation noise. In this paper, we propose U-4DGS, a framework integrating a Probabilistic Deformation Network and a Double Rasterization pipeline. This architecture renders pixel-aligned uncertainty maps that act as an adaptive gradient modulator, automatically attenuating artifacts from unreliable observations. Furthermore, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations, which leverage the learned uncertainty to selectively propagate spatial-temporal validity. Extensive experiments on ZJU-MoCap and OcMotion demonstrate that U-4DGS achieves SOTA rendering fidelity and robustness.
1 Introduction
@@
High-fidelity reconstruction of dynamic humans from monocular videos stands as a foundational technology for immersive applications, spanning telepresence, metaverse interaction, sports analysis, and virtual reality [tian2023recovering, yang2025innovative]. Recent years have witnessed remarkable progress in this domain, enabling the synthesis of photorealistic digital avatars from monocular inputs [weng2022humannerf, wen2024gomavatar]. However, these successes hinge on the idealized assumption of full visibility and isolation from the background. This assumption rarely holds in unconstrained in-the-wild scenarios, where subjects inevitably interact with their surroundings, such as sitting on chairs, navigating behind obstacles, or being partially obscured by environmental objects. When standard pipelines are applied to occluded scenarios, their performance degrades catastrophically [xiang2023rendering, sun2024occfusion]. As illustrated in Fig. 1(e), the model erroneously interprets the occluder as part of the human geometry, manifesting as incomplete geometry or severe texture artifacts. Consequently, enabling robust human rendering from occluded observations is imperative to bridge the gap between laboratory prototypes and real-world applications.
@@
To address these specific challenges, pioneering NeRF-based approaches, such as OccNeRF [xiang2023rendering] and Wild2Avatar [xiang2025rendering], explore the task of rendering humans from object-occluded monocular videos. By incorporating geometry priors or decoupling scene components, these methods demonstrate the feasibility of recovering occluded surfaces. However, despite their visual fidelity, they inherit the intrinsic computational bottlenecks of implicit neural representations. Specifically, the reliance on expensive volumetric rendering incurs prohibitive training and inference costs, effectively precluding their deployment in real-time applications.
@@
Recently, the advent of 3D Gaussian Splatting (3DGS) [kerbl20233d] shifted the paradigm, offering real-time rendering capabilities [fei20243d, feng2025flashgs]. To adapt 3DGS for occluded human rendering, concurrent works incorporate external priors to compensate for missing information. Methods relying on geometric heuristics enforce rigid rules, such as left-right symmetry [jiang2025symgaussian] or feature aggregation [ye2025occgaussian]. However, these hand-crafted spatial priors typically lack flexibility in modeling diverse human appearances. As shown in Fig. 1(f), the symmetry prior fails to recover the chest logo, erroneously copying unadorned features from the opposite side. Alternatively, methods leveraging generative priors utilize 2D diffusion models to hallucinate invisible regions [sun2024occfusion, lee2024guess]. Yet, the inherent stochasticity of frame-by-frame generation often induces severe temporal flickering. This is evident in Fig. 1(g), where the shirt color drifts inconsistently across frames, violating physical consistency. Crucially, these methods rely on imperfect external priors to infer missing content, neglecting the intrinsic reliability of the input data, where valid cues are embedded across the temporal sequence.
@@
To bridge this gap, our insight stems from the video’s intrinsic temporal redundancy: since occlusions are transient, parts obscured in one frame are likely visible in others. This motivates adopting a 4D representation to consolidate fragmented observations into a unified canonical space. However, naive 4D Gaussian Splatting faces a supervision dilemma [wu20244d, li2024st]. Indiscriminately treating all pixels as valid constraints forces the deformation field to establish erroneous correspondences, compelling the human geometry to adhere to the occluder’s surface texture. Conversely, strictly limiting supervision to visible regions leaves the occluded parts mathematically unconstrained; without gradient feedback, the geometry in these blind spots fails to densify or maintain coherence, resulting in incomplete structures. Thus, the reconstruction process demands a dynamic discrimination mechanism to assess the validity of the observation, differentiating between visible human details and misleading observational noise (i.e., occluded regions).
@@
Building upon this formulation, we present U-4DGS, a robust uncertainty-aware framework for monocular occluded human rendering. To parameterize the heteroscedastic noise model, we propose a Probabilistic Deformation Network. Beyond standard geometric warping, this module explicitly predicts the per-primitive aleatoric uncertainty, serving as the confidence metric for our MAP objective. To enable the adaptive gradient modulation, we introduce a Double Rasterization pipeline. By rendering pixel-aligned uncertainty maps alongside color, this mechanism enforces the inverse weighting scheme, ensuring that gradients from occluded regions are effectively attenuated. Finally, to prevent geometric drift in areas lacking visual cues (i.e., occluded regions), we design Confidence-Aware Regularizations. These constraints leverage the learned uncertainty to selectively propagate validity from confident regions to unreliable areas, ensuring coherent completion. Consequently, as illustrated in Fig. 1(a), U-4DGS achieves high-fidelity rendering even under severe occlusion while maintaining training efficiency.
@@
We propose U-4DGS, a framework integrating a Probabilistic Deformation Network with a Double Rasterization pipeline. This design renders pixel-aligned uncertainty maps that act as adaptive gradient modulators, effectively shielding the canonical geometry from corruption.
@@ 2 Related Works
3D Human Avatar Reconstruction. Traditional human reconstruction methods typically rely on dense camera arrays or depth sensors, limiting their application in in-the-wild scenarios [collet2015high, su2020robustfusion, huang2020arch, yu2021function4d]. To address this, NeRF-based methods have enabled high-fidelity reconstruction from monocular inputs by conditioning radiance fields on SMPL priors [yu2023monohuman, te2022neural, li2024ghunerf]. However, the reliance on volumetric ray-marching incurs prohibitive computational costs, precluding real-time applicability. Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has introduced a paradigm shift with its real-time rasterization capabilities. State-of-the-art methods [qian20243dgs, kocabas2024hugs, hu2024gauhuman, lei2024gart, shao2024splattingavatar, zhao2025surfel] adapt this representation to dynamic humans by representing the body in a canonical space and transforming it via Linear Blend Skinning. Critically, while performant under full visibility, these standard pipelines implicitly assume valid supervision for all pixels. Thus, they lack mechanisms to distinguish the human subject from occlusions, leading to catastrophic degradation where occluder artifacts are erroneously fused into the avatar geometry.
@@
Occlusion-Aware Human Rendering. Existing methodologies generally fall into three paradigms. Early strategies focused on scene decoupling, employing visibility priors to isolate the subject from obstacles [xiang2025rendering, xiang2023rendering]; however, these implicit frameworks often suffer from high computational overhead. To synthesize unobserved content, recent approaches leverage generative priors via pre-trained 2D diffusion models [sun2024occfusion, lee2024guess, fan2026inpainthuman]. Yet, the inherent stochasticity of such processes frequently induces identity drift and temporal flickering. Alternatively, other 3DGS-based methods employ geometric heuristics such as symmetry constraints [ye2025occgaussian, jiang2025symgaussian], which often falter under asymmetric clothing or motions. Distinctively, U-4DGS eschews unstable hallucinations and rigid heuristics. Instead, we formulate the reconstruction as an MAP estimation problem. By strictly aggregating valid temporal information via an adaptive gradient modulation mechanism, we ensure reconstruction that is both physically consistent and faithful to the subject.
@@
Dynamic Reconstruction and Uncertainty Modeling. Recent strides in dynamic Gaussian Splatting have achieved high-fidelity 4D reconstruction via deformation fields [yang2024deformable, wu20244d, luiten2024dynamic, bae2024per, zhang2025mega]. However, these methods predicate their success on a deterministic correspondence between canonical and observation spaces. While valid for clean sequences, this premise disintegrates under occlusion, where the absence of valid motion cues causes the solver to overfit to occluders [guo2025uncertainty]. To address this, uncertainty modeling has been explored to enhance robustness against sparse views or transient noise [sunderhauf2023density, goli2024bayes, ren2024nerf]. Notably, concurrent work like USPLAT4D [guo2025uncertainty] employs an uncertainty-aware motion graph for general dynamic scenes. Yet, treating uncertainty solely as a smoothing weight or edge attribute proves insufficient for articulated human reconstruction, where structural integrity is paramount. In contrast, U-4DGS elevates uncertainty to a central arbitration role. Beyond simple loss attenuation, we leverage learned confidence to actively modulate optimization dynamics: shielding canonical geometry from misleading gradients while simultaneously triggering confidence-aware regularizations to enforce physical plausibility within regions lacking reliable visual cues.
3 Preliminaries
3.1 3D Gaussian Splatting
@@ 3 Preliminaries
3D Gaussian Splatting (3DGS) [kerbl20233d] explicitly models the scene using a set of anisotropic 3D Gaussians. Each primitive is parameterized by a mean , covariance , opacity , and view-dependent color (via Spherical Harmonics). To enforce positive semi-definiteness during differentiable optimization, the covariance is decomposed into a rotation quaternion and a scaling vector :
@@
3.2 Articulated Human Gaussians
Following recent works [hu2024gauhuman, qian20243dgs], we leverage the SMPL model [loper2015smpl] to establish a canonical coordinate system. SMPL provides a skinned vertex mesh parametrized by pose and shape . To model dynamic humans, we instantiate 3D Gaussians within this canonical T-pose space. For a specific frame with pose , any canonical point is transformed to the observation space via Linear Blend Skinning:
@@ 3.2 Articulated Human Gaussians
3.3 Problem Formulation: A Probabilistic Perspective
@@
4 Methodology
4.1 Overview
Building upon the probabilistic formulation derived in Sec. 3, we propose U-4DGS, a framework designed to realize the MAP estimation for robust occluded human rendering. U-4DGS explicitly models the heteroscedastic nature of observation noise, enabling an autonomous arbitration mechanism between valid supervision and occlusion artifacts. As illustrated in Fig. 2, the pipeline consists of three integral components: (1) Probabilistic Deformation (Sec. 4.2): We introduce a network that serves a dual role: it models the temporal geometry evolution via deformation offsets and parameterizes the input-dependent aleatoric uncertainty for each Gaussian, explicitly quantifying the confidence of the temporal correspondence. (2) Double Rasterization (Sec. 4.3): To propagate this 3D confidence to the 2D observation space, we employ a dual-branch rasterizer. This module renders the deformed geometry into a photometric image while simultaneously accumulating per-primitive uncertainty into a pixel-aligned Uncertainty Map . (3) Optimization via MAP Estimation (Sec. 4.4): These outputs drive our MAP-based objective. Here, functions as an adaptive gradient modulator within the probabilistic photometric loss, attenuating gradients from unreliable observations. Simultaneously, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations that leverage the learned uncertainty to impose spatial-temporal priors.
4.2 Probabilistic Deformation Network
To instantiate the MAP framework, we require a mechanism that not only recovers time-varying geometry but also quantifies the reliability of each correspondence. Unlike standard deterministic fields [yang2024deformable, hu2024gauhuman] that indiscriminately fit all observations, we propose a Probabilistic Deformation Network . This module serves a dual purpose: (1) it models high-frequency non-rigid residuals beyond the SMPL prior; (2) it explicitly predicts the aleatoric uncertainty , providing the input-dependent noise scale required for our probabilistic objective.
@@ 4.2 Probabilistic Deformation Network
4.3 Double Rasterization
To bridge the gap between the 3D aleatoric uncertainty and the 2D observation space required for our MAP objective, we propose a Double Rasterization scheme. Since occlusion is an inherently view-dependent phenomenon, the reliability of a pixel is determined by the accumulated uncertainty along its corresponding optical ray. Therefore, we perform two pixel-aligned rasterization passes in parallel, generating both the photometric appearance and the uncertainty map .

### v2

@@ Abstract.
High-fidelity rendering of dynamic humans from monocular videos typically degrades catastrophically under occlusions. Existing solutions incorporate external priors—either hallucinating missing content via generative models, which induces severe temporal flickering, or imposing rigid geometric heuristics that fail to capture diverse appearances. To this end, we reformulate the task as a Maximum A Posteriori estimation problem under heteroscedastic observation noise. In this paper, we propose U-4DGS, a framework integrating a Probabilistic Deformation Network and a Joint Rasterization pipeline. This architecture renders pixel-aligned uncertainty maps that act as an adaptive gradient modulator, automatically attenuating artifacts from unreliable observations. Furthermore, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations, which leverage the learned uncertainty to selectively propagate spatial-temporal validity. Extensive experiments on the ZJU-MoCap and OcMotion datasets demonstrate that U-4DGS achieves state-of-the-art rendering fidelity and robustness.
1. Introduction
@@
High-fidelity reconstruction of dynamic humans from monocular videos stands as a foundational technology for immersive applications, spanning telepresence, metaverse interaction, sports analysis, and virtual reality (Tian et al., 2023; Yang et al., 2025). Recent years have witnessed remarkable progress in this domain, enabling the synthesis of photorealistic digital avatars from monocular inputs (Weng et al., 2022; Wen et al., 2024). However, these successes hinge on the idealized assumption of full visibility and isolation from the background. This assumption rarely holds in unconstrained in-the-wild scenarios, where subjects inevitably interact with their surroundings, such as sitting on chairs, navigating behind obstacles, or being partially obscured by environmental objects. When standard pipelines are applied to occluded scenarios, their performance degrades catastrophically (Xiang et al., 2023; Sun et al., 2024). As illustrated in Fig. 1(e), the model erroneously interprets the occluder as part of the human geometry, manifesting as incomplete geometry or severe texture artifacts. Consequently, enabling robust human rendering from occluded observations is imperative to bridge the gap between laboratory prototypes and real-world applications.
@@
To address these specific challenges, pioneering NeRF-based approaches, such as OccNeRF (Xiang et al., 2023) and Wild2Avatar (Xiang et al., 2025), explore the task of rendering humans from object-occluded monocular videos. By incorporating geometry priors or decoupling scene components, these methods demonstrate the feasibility of recovering occluded surfaces. However, despite their visual fidelity, they inherit the intrinsic computational bottlenecks of implicit neural representations. Specifically, the reliance on expensive volumetric rendering incurs prohibitive training and inference costs, effectively precluding their deployment in real-time applications.
@@
Recently, the advent of 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) shifted the paradigm, offering real-time rendering capabilities (Fei et al., 2024; Feng et al., 2025). To adapt 3DGS for occluded human rendering, concurrent works incorporate external priors to compensate for missing information. Methods relying on geometric heuristics enforce rigid rules, such as left-right symmetry (Jiang et al., 2025) or feature aggregation (Ye et al., 2025). However, these hand-crafted spatial priors typically lack flexibility in modeling diverse human appearances. As shown in Fig. 1(f), the symmetry prior fails to recover the chest logo, erroneously copying unadorned features from the opposite side. Alternatively, methods leveraging generative priors utilize 2D diffusion models to hallucinate invisible regions (Sun et al., 2024; Lee et al., 2024). Yet, the inherent stochasticity of frame-by-frame generation often induces severe temporal flickering. This is evident in Fig. 1(g), where the shirt color drifts inconsistently across frames, violating physical consistency. Crucially, these methods rely on imperfect external priors to infer missing content, neglecting the intrinsic reliability of the input data, where valid cues are embedded across the temporal sequence.
@@
To bridge this gap, our insight stems from the video’s intrinsic temporal redundancy: since occlusions are transient, parts obscured in one frame are likely visible in others. This motivates adopting a 4D representation to consolidate fragmented observations into a unified canonical space. However, naive 4D Gaussian Splatting faces a supervision dilemma (Wu et al., 2024; Li et al., 2024b). Indiscriminately treating all pixels as valid constraints forces the deformation field to establish erroneous correspondences, compelling the human geometry to adhere to the occluder’s surface texture. Conversely, a seemingly straightforward alternative is to apply hard visibility masks to exclude occluded regions from the optimization. Yet, strictly limiting supervision to visible regions leaves the occluded parts mathematically unconstrained; without gradient feedback, the geometry in these blind spots fails to densify or maintain coherence, resulting in incomplete structures. Thus, rather than relying on brittle hard masks, the reconstruction process demands a dynamic discrimination mechanism to assess the validity of the observation, differentiating between visible human details and misleading observational noise (i.e., occluded regions).
@@
Building upon this formulation, we present U-4DGS, a robust uncertainty-aware framework for monocular occluded human rendering. To parameterize the heteroscedastic noise model, we propose a Probabilistic Deformation Network. Beyond standard geometric warping, this module explicitly predicts the per-primitive aleatoric uncertainty, serving as the confidence metric for our MAP objective. To enable the adaptive gradient modulation, we introduce a Joint Rasterization pipeline. By rendering pixel-aligned uncertainty maps alongside color, this mechanism enforces the inverse weighting scheme, ensuring that gradients from occluded regions are effectively attenuated. Finally, to prevent geometric drift in areas lacking visual cues (i.e., occluded regions), we design Confidence-Aware Regularizations. These constraints leverage the learned uncertainty to selectively propagate validity from confident regions to unreliable areas, ensuring coherent completion. As illustrated in Fig. 1(a), U-4DGS achieves high-fidelity rendering even under severe occlusion while maintaining training efficiency.
@@
We propose U-4DGS, a framework integrating a Probabilistic Deformation Network with a Joint Rasterization pipeline. This design renders pixel-aligned uncertainty maps that act as adaptive gradient modulators, effectively shielding the canonical geometry from corruption.
@@ 2. Related Works
3D Human Avatar Reconstruction. Traditional human reconstruction methods typically rely on dense camera arrays or depth sensors, limiting their application in in-the-wild scenarios (Collet et al., 2015; Su et al., 2020; Huang et al., 2020; Yu et al., 2021). To address this, NeRF-based methods have enabled high-fidelity reconstruction from monocular inputs by conditioning radiance fields on SMPL priors (Yu et al., 2023; Te et al., 2022; Li et al., 2024a). However, the reliance on volumetric ray-marching incurs prohibitive computational costs, precluding real-time applicability. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has introduced a paradigm shift with its real-time rasterization capabilities. Building upon the skeleton-conditioned canonical deformation paradigm pioneered by early implicit frameworks like Animatable NeRF (Peng et al., 2021a), SOTA methods (Qian et al., 2024; Kocabas et al., 2024; Hu et al., 2024b; Lei et al., 2024; Shao et al., 2024; Zhao et al., 2025) adapt this representation to dynamic humans by representing the body in a canonical space and transforming it via Linear Blend Skinning. Critically, while performant under full visibility, these standard pipelines implicitly assume valid supervision for all pixels. Thus, they lack mechanisms to distinguish the human subject from occlusions, leading to catastrophic degradation where occluder artifacts are erroneously fused into the avatar geometry.
@@
Occlusion-Aware Human Rendering. Existing methodologies generally fall into three paradigms. Early strategies focused on scene decoupling, employing visibility priors to isolate the subject from obstacles (Xiang et al., 2025, 2023); however, these implicit frameworks often suffer from high computational overhead. To synthesize unobserved content, recent approaches leverage generative priors via pre-trained 2D diffusion models (Sun et al., 2024; Lee et al., 2024; Fan et al., 2026). Yet, the inherent stochasticity of such processes frequently induces identity drift and temporal flickering. Alternatively, other 3DGS-based methods employ geometric heuristics such as symmetry constraints (Ye et al., 2025; Jiang et al., 2025), which often falter under asymmetric clothing or motions. Distinctively, U-4DGS eschews unstable hallucinations and rigid heuristics. Instead, we formulate the reconstruction as an MAP estimation problem. By strictly aggregating valid temporal information via an adaptive gradient modulation mechanism, we ensure reconstruction that is both physically consistent and faithful to the subject.
@@
Dynamic Reconstruction and Uncertainty Modeling. Recent strides in dynamic Gaussian Splatting have achieved high-fidelity 4D reconstruction via deformation fields (Yang et al., 2024; Wu et al., 2024; Luiten et al., 2024; Bae et al., 2024; Zhang et al., 2025). However, these methods predicate their success on a deterministic correspondence between canonical and observation spaces. While valid for clean sequences, this premise disintegrates under occlusion, where the absence of valid motion cues causes the solver to overfit to occluders (Guo et al., 2025). To address this, uncertainty modeling has been explored to enhance robustness against sparse views or transient noise (Sünderhauf et al., 2023; Goli et al., 2024; Ren et al., 2024). Grounded in Bayesian deep learning for aleatoric uncertainty (Kendall and Gal, 2017), prior neural rendering methods have successfully mitigated transient distractors utilizing strategies such as per-image uncertainty embeddings (Martin-Brualla et al., 2021), robust loss formulations (Sabour et al., 2023), or direct volume density parameterization (Lee et al., 2025). Concurrent work like USPLAT4D (Guo et al., 2025) employs an uncertainty-aware motion graph for general dynamic scenes. Yet, treating uncertainty solely as a smoothing weight or edge attribute proves insufficient for articulated human reconstruction, where structural integrity is paramount. In contrast, U-4DGS elevates uncertainty to a central arbitration role. Beyond simple loss attenuation, we leverage learned confidence to actively modulate optimization dynamics: shielding canonical geometry from misleading gradients while simultaneously triggering confidence-aware regularizations to enforce physical plausibility within regions lacking reliable visual cues.
3. Preliminaries
3.1. 3D Gaussian Splatting
@@ 3. Preliminaries
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) explicitly models the scene using a set of anisotropic 3D Gaussians. Each primitive is parameterized by a mean , covariance , opacity , and view-dependent color (via Spherical Harmonics). To enforce positive semi-definiteness during differentiable optimization, the covariance is decomposed into a rotation quaternion and a scaling vector :
@@
3.2. Articulated Human Gaussians
Following recent works (Hu et al., 2024b; Qian et al., 2024), we leverage the SMPL model (Loper et al., 2015) to establish a canonical coordinate system. SMPL provides a skinned vertex mesh parametrized by pose and shape . To model dynamic humans, we instantiate 3D Gaussians within this canonical T-pose space. For a specific frame with pose , any canonical point is transformed to the observation space via Linear Blend Skinning:
@@ 3.2. Articulated Human Gaussians
3.3. Problem Formulation
@@
4. Methodology
4.1. Overview
Building upon the probabilistic formulation derived in Sec. 3, we propose U-4DGS, a framework designed to realize the MAP estimation for robust occluded human rendering. U-4DGS explicitly models the heteroscedastic nature of observation noise, enabling an autonomous arbitration mechanism between valid supervision and occlusion artifacts. As illustrated in Fig. 2, the pipeline consists of three integral components: (1) Probabilistic Deformation (Sec. 4.2): We introduce a network that serves a dual role: it models the temporal geometry evolution via deformation offsets and parameterizes the input-dependent aleatoric uncertainty for each Gaussian, explicitly quantifying the confidence of the temporal correspondence. (2) Joint Rasterization (Sec. 4.3): To propagate this 3D confidence to the 2D observation space, we employ a dual-branch rasterizer. This module renders the deformed geometry into a photometric image while simultaneously accumulating per-primitive uncertainty into a pixel-aligned Uncertainty Map . (3) Optimization via MAP Estimation (Sec. 4.4): These outputs drive our MAP-based objective. Here, functions as an adaptive gradient modulator within the probabilistic photometric loss, attenuating gradients from unreliable observations. Simultaneously, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations that leverage the learned uncertainty to impose spatial-temporal priors.
4.2. Probabilistic Deformation Network
To instantiate the MAP framework, we require a mechanism that not only recovers time-varying geometry but also quantifies the reliability of each correspondence. Unlike standard deterministic fields (Yang et al., 2024; Hu et al., 2024b) that indiscriminately fit all observations, we propose a Probabilistic Deformation Network . This module serves a dual purpose: (1) it models high-frequency non-rigid residuals beyond the SMPL prior; (2) it explicitly predicts the aleatoric uncertainty , providing the input-dependent noise scale required for our probabilistic objective.
@@ 4.2. Probabilistic Deformation Network
[…]
| Standard Human Rendering | |||||||||||||
| HumanNeRF (Weng et al., 2022) | - | 20.67 | 0.9509 | - | - | - | - | ||||||
| GaussianAvatar (Hu et al., 2024a) | - | 18.01 | 0.9512 | 60.33 | - | - | - | ||||||
| GauHuman (Hu et al., 2024b) | - | 21.55 | 0.9430 | 55.88 | 15.09 | 0.8525 | 107.1 | ||||||
| Occlusion-Aware Approaches | |||||||||||||
| OccNeRF (Xiang et al., 2023) | Scene Decoupling | 22.40 | 0.9562 | 43.01 | 15.71 | 0.8523 | 82.90 | ||||||
| Wild2Avatar (Xiang et al., 2025) | - | - | - | 14.09 | 0.8484 | 93.31 | |||||||
| OccGaussian (Ye et al., 2025) | Geometric Heuristics | 23.29 | 0.9482 | 41.93 | - | - | - | ||||||
| SymGaussian (Jiang et al., 2025) | 23.22 | 0.9535 | 39.02 | - | - | - | |||||||
| GTU (Lee et al., 2024) | Generative Priors | 22.89 | 0.9503 | 40.78 | 15.83 | 0.8437 | 83.46 | ||||||
| OccFusion (Sun et al., 2024) | 23.96 | 0.9548 | 32.34 | 18.28 | 0.8875 | 82.42 | |||||||
| U-4DGS (Ours) | 24.62 | 0.9606 | 31.72 | 20.11 | 0.9030 | 79.17 | |||||||
4.3. Joint Rasterization
To bridge the gap between the 3D aleatoric uncertainty and the 2D observation space required for our MAP objective, we propose a Joint Rasterization scheme. Since occlusion is an inherently view-dependent phenomenon, the reliability of a pixel is determined by the accumulated uncertainty along its corresponding optical ray. Therefore, we perform two pixel-aligned rasterization passes in parallel, generating both the photometric appearance and the uncertainty map .